# Weichuang Shao Iman Yi Liao Tomas Henrique Bode Maul Tissa Chandesa

### v1

@@ 1 Introduction
Audio classification is a machine learning task that aims to identify semantic categories from audio signals. Pre-trained models such as SSAST [16], EAT [6], HuBERT [19], Qwen-Audio [8], and Audio Flamingo [25], have achieved state-of-the-art performance on benchmark datasets including SpeechCommands V1 (SC1) [42], SpeechCommands V2 (SC2) [42], AudioMNIST (AM) [4], VocalSound (VS) [17, 15], and CochlScene (CS) [21]. However, these models rely heavily on pre-training to reach such performance levels, which imposes several practical limitations. Specifically, reusing pre-trained weights from these models requires a fixed sample rate (16 kHz for SSAST, EAT, HuBERT, and Qwen-audio, and 44.1 kHz for Audio Flamingo) and specific audio length (1 s, 5 s, or 10 s). These constraints restrict model reusability and make it difficult to apply the models to new datasets with different sample rates or audio lengths.
Although resampling techniques can be applied to meet the required rates, such preprocessing often introduces information loss when downsampling or artificial data when upsampling. For instance, [29] reported that the main frequency of bird sounds is below 12 kHz, implying that, based on the information theory [39] (Nyquist theorem), the sample rate should exceed 24 kHz to preserve information. However, 24 kHz is not a standard rate supported by most existing pre-trained models. Furthermore, as audio recording hardware continues to evolve, producing higher-quality data at higher sample rates, a fixed sample rate methodology cannot take advantage of such an improvement in the recording device.
The same limitation applies to audio length. SSAST and EAT, for example, support only predefined durations of 1, 5, or 10 seconds [16, 6], while Audio Flamingo allows a maximum of 33.25 seconds [25]. When the input audio length or sample rate does not match the model’s configuration, pre-training must be repeated or the audio must be forcibly adjusted, both of which are computationally expensive.
@@
    Dual CLS + TAL tokens: a novel token pairing that captures bidirectional contextual information within the Transformer encoder.
  4. 4.
@@ 2.1 Transformer-based Audio Models
The transformer architecture, initially developed for language processing, has been effectively adapted for vision and audio tasks. Vision Transformer (ViT) [11] and Distilled ViT (DeiT) [35] operate by dividing images into fixed-size patches (e.g., $16\times 16$ or $32\times 32$), a concept later transferred to the audio domain, such as Audio Spectrogram Transformer (AST) [14]. AST represents audio as Mel-spectrograms [36, 20], turning temporal signals into two-dimensional time–frequency grids before applying patch-based Transformer processing with fixed-size patches ($16\times 16$). While AST reuses ImageNet-trained weights [9, 14], SSAST [16] processes pre-training on AudioSet [13] and LibriSpeech [30] rather than reusing ImageNet-trained weights for extending AST.
Models such as Wav2Vec 2.0 [3] and HuBERT [10] process raw wavforms using 1D CNN-Transformer architectures. Wav2Vec 2.0 is composed of seven convolutional layers equipped with LayerNorm and GELU activations, and its large-scale pre-training on LibriSpeech [30] and Libri-Light [22] reportedly required about 7.5 days on 128 GPUs. HuBERT follows a similar CNN-Transformer architecture to Wav2Vec 2.0 but extends it with a BERT-style [10] masked prediction mechanism that integrates a waveform encoder. Like Wav2Vec 2.0, AMAuT adopts a 1D CNN-Transformer structure but differs fundamentally: AMAuT trains from scratch without pre-training, uses 1D CNNs on Mel-spectrogram horizontal tokens, and replaces LayerNorm + GELU with BatchNorm + ReLU. These changes improve both efficiency and accuracy (99.82% vs. 99.4% on AudioMNIST) between Wav2Vec 2.0 and AMAuT.
The spectrogram transformers [43] contributed a Tempo-Frequency sequential (TFS) attention for separating mel-spectrogram outputs within frequency direction instead of $16\times 16$ regions in AST [14] that are similar to the horizontal tokenization of AMAuT (see subsection 3.3). However, [43] inherits the methodology of AST and ViT, and then [43] employs a different model architecture, as it incorporates 2D CNN layers.
@@ 2.2 Test-Time Domain Adaptation
Test-time domain adaptation (TTDA) only uses the test set for performance improvement by adapting (partial) model parameters of a trained model from a source domain to a target domain [28]. In AMAuT, this is achieved through a composite loss function $\mathcal{L}_{TTDA}$ that combines three complementary objectives: Nuclear-Norm Maximization, Entropy Minimization, and Generalized Entropy (see Eq.7, Eq.8, and Eq.9 in Appendix A). These components have been applied separately in previous studies, such as CoNMix [26] for test-time domain adaptation, TENT [41] for entropy-based test-time batch adaptation, and SGEM [23] for Automatic Speech Recognition, but AMAuT unifies them for the first time. This integration allows efficient test-time domain adaptation and improves robustness under moderate distribution shifts without additional supervision.
@@ 3.2 Multiview Augmentation
  • Training (four augmentations): (1) random left or right time shift (0-17%), (2) Gaussian noise (noise-to-audio ratio = 0.015, and regenerate noise each time), and (3-4) random clips from two background noises (“dude miaowing” and “pick noise” from SpeechCommands [42]) with an SNR of 50 dB.
@@ 3.5 Vertical Embedding
AMAuT introduces two learnable tokens: the class token (CLS) and the tail token (TAL) (see Eq. 3 and Figure 3). Unlike conventional designs such as DeiT [35], where CLS and distillation tokens are placed adjacently, AMAuT positions CLS and TAL at opposite ends of the token sequence. This design assumes that tokens at different spatial positions may encode distinct contextual information during transformer processing.
@@
This component follows the Transformer [38] architecture and comprises multiple transformer blocks. Two important configurations are adopted:
@@
    The dropout rate for both multi-head attention and MLP is set to 0%, consistent with models such as SSAST [16], AST [14], and CoNMix [26].
@@ 3.7 Test-time Adaptation / Augmentation ($TTA^{2}$)
This subsection investigates an uncommon application of Test-time Domain Adaptation (TTDA): improving accuracy when training and test distributions are similar. Typically, TTDA is employed to address distribution shifts, but we explore its utility even when such a shift is minimal [28].
TTDA enhances performance by adapting a trained model to the test data distribution without requiring labels [28]. During TTDA, all test features are visible, but labels remain hidden. The model updates its internal parameters to better align with the target domain before final prediction.
@@
where $\alpha$, $\beta$, and $\gamma$ are scaling coefficients (see Table11 in Appedix B).
@@
To further enhance inference robustness, AMAuT adopts test-time augmentation (TTAu) [24]. Unlike TTDA, TTAu does not update model parameters; instead, it refines predictions by averaging outputs from multiple augmented versions of the same input. Three TTAu strategies are evaluated: Augmentation-driven refinement (Aug), Multi-training-based refinement (Mlt), and Hybrid refinement (Hyb) (see Eq.10, Eq.11, and Eq.12 in Appendix A, separately).
@@ 3.8 Learning Rate and Optimization
where $e$ denotes the epoch number, $10\leq\lambda\leq 40$ controls the rate of change, $200\geq\eta\geq 40$ defines the cardinality. See more details of hyperparameter settings from Table 12 in Appendix B.
@@ 4 Experiments and Results
AMAuT integrates several architectural and training strategies, including augmentation-driven multiview learning, conv1 + conv7 + conv1 CNN bottlenecks, flexible handling of arbitrary sample rates and audio lengths, CLS+TAL learnable positional embeddings, and $TTA^{2}$. Collectively, these components enable AMAuT to achieve strong performance when trained entirely from scratch, without relying on large-scale pre-training.
@@ 4.1 Datasets
To evaluate the proposed AMAuT framework, this study employed five publicly available benchmark datasets: AudioMNIST (AM) [4], SpeechCommands V1 (SC1) [42], SpeechCommands V2 (SC2) [42], VocalSound (VS) [17, 15], and CochlScene (CS) [21].
@@ 4.2 Ablation Studies
Table 3: Comparison of test accuracies between sequential augmentation and augmentation-driven multiview learning on CochlScene (CS) and SpeechCommands V1 (SC1) before applying $TTA^{2}$. “1-Aug” and “4-Aug” indicate the use of one or four augmented views, respectively. The arrows ($\uparrow$, $\downarrow$) denote performance improvement or degradation relative to the baseline.
@@
In many existing CNN-based audio classification architectures, relatively small convolutional kernels are commonly used. For example, HuBERT [19] and Wav2Vec 2.0 [3] employ bottleneck structures such as conv2 + conv3, while other hybrid CNN-Transformer architectures like ENACT-Heart [18] and CNN-TE [5] use kernel sizes of 3, 2, and 1. These designs typically include only one or two large convolutional layers, such as conv14 or conv16, at the head or tail of the CNN stack, rather than throughout its body.
@@
Table 4: Performance comparison between conv3 and conv7 kernel structures on CochlScene (CS) and SpeechCommands V1 (SC1) before applying $(TTA)^{2}$, evaluated on a single RTX 4090 GPU. “Length” indicates audio length in seconds, and “1-Aug” / “4-Aug” denote the number of augmented views used during testing.
@@
CLS+TAL Tokens
In AMAuT, the CLS (class) and TAL (tail) tokens are positioned at opposite ends of the token sequence, rather than adjacent to each other as in conventional Transformer designs. This placement allows the model to capture complementary contextual information from both the beginning and end of the sequence, enhancing the overall representational diversity during attention processing.
Table 5: Comparing different positional embedding strategies on CochlScene (CS) and VocalSound (VS) before applying $TTA^{2}$. The tested configurations include CLS+TAL (proposed), CLS+Distil [35], and CLS only [11]. ’Length’ denotes audio length in seconds.
Set Length Method Accuracy (%)
CS 10 s CLS+TAL 68.87
CS 10 s CLS+Distil 69.49
CS 10 s CLS 67.13
VS 12 s CLS+TAL 91.06
VS 12 s CLS+Distil 90.61
VS 12 s CLS 90.56
The comparative results summarized in Table 5 demonstrate that the CLS+TAL positional embedding mechanism introduced in this study serves as an alternative to the commonly used CLS+Distil embedding configuration proposed in DeiT [35].
@@
Table 6: Test accuracy improvement (%) from different refinement strategies, including Augmentation (Aug), Multi-training (Mlt), Hybrid (Hyb), and Test-time Domain Adaptation (TTDA), applied to AMAuT. For AudioMNIST (AM), results are reported over five folds (AM0–AM4).
@@
Table 7: Adaptation time comparison for different $TTA^{2}$ methods in AMAuT using a single RTX 4090 GPU. Time is shown in seconds (s) or minutes (m) per method. “TTDA-Epoch” denotes time costing and the number of epochs used for performing TTDA.
@@
4.3 Comparative Performance
Table 8: Comparison of prediction accuracy (%) between the proposed AMAuT and state-of-the-art pre-trained models on five benchmark datasets. Performance values for baseline models are obtained from their respective original publications.
Algorithm Set Accuracy(%)
AMAuT SC1 96.31
SSAST [16] SC1 96.20
AMAuT SC2 96.21
EAT [6] SC2 98.30
AST [14] SC2 98.11
PDC [7] SC2 97.8
AMAuT AM 99.82
HuBERT [19, 27] AM 99.95
XLS-R [2, 27] AM 99.86
Wav2Vec2 [3, 27] AM 99.42
AMAuT VS 92.26
Qwen-Audio [8] VS 92.89
EfficientNet [17, 34] VS 90.5
AMAuT CS 72.58
Audio Flamingo [25] CS 83.0
Qwen-Audio [8] CS 79.5
Table 8 compares AMAuT with several state-of-the-art pre-trained models across the five datasets. AMAuT achieved 96.31% on SC1 and 99.82% on AM, closely matching or exceeding prior results, while maintaining minimal computational overhead.

### v2

@@ 1 Introduction
Audio classification is a machine learning task that aims to identify semantic categories from audio signals. Pre-trained models such as SSAST [gong2022ssast], EAT [chen2024eat], HuBERT [hsu2021hubert], Qwen-Audio [chu2023qwen], and Audio Flamingo [kong2024audio], have achieved state-of-the-art performance on benchmark datasets including SpeechCommands V1 (SC1) & V2 (SC2) [warden2018speech], AudioMNIST (AM) [audiomnist2023], VocalSound (VS) [gong_vocalsound], and CochlScene (CS) [jeong2022cochlscene]. However, these models rely heavily on pre-training to reach such performance levels, which imposes several practical limitations. Specifically, reusing pre-trained weights from these models requires a fixed sample rate (16 kHz for SSAST, EAT, HuBERT, and Qwen-audio, and 44.1 kHz for Audio Flamingo) and specific audio length (1 s, 5 s, or 10 s). These constraints restrict model reusability and make it difficult to apply the models to new datasets with different sample rates or audio lengths.
Although resampling techniques can be applied to meet the required rates, such preprocessing often introduces information loss when downsampling or artificial data when upsampling. For instance, [loo2025temporal] reported that the main frequency of bird sounds is below 12 kHz, implying that, based on the information theory [volkenstein2009entropy] (Nyquist theorem), the sample rate should exceed 24 kHz to preserve information. However, 24 kHz is not a standard rate supported by most existing pre-trained models. Furthermore, as audio recording hardware continues to evolve, producing higher-quality data at higher sample rates, a fixed sample rate methodology cannot take advantage of such an improvement in the recording device.
The same limitation applies to audio length. SSAST and EAT, for example, support only predefined durations of 1, 5, or 10 seconds [gong2022ssast, chen2024eat], while Audio Flamingo allows a maximum of 33.25 seconds [kong2024audio]. When the input audio length or sample rate does not match the model’s configuration, pre-training must be repeated or the audio must be forcibly adjusted, both of which are computationally expensive.
@@ 2.1 Transformer-based Audio Models
The transformer architecture, initially developed for language processing, has been effectively adapted for vision and audio tasks. Vision Transformer (ViT) [dosovitskiy2020image] and Distilled ViT (DeiT) [touvron2021training] operate by dividing images into fixed-size patches (e.g., $16\times 16$ or $32\times 32$), a concept later transferred to the audio domain, such as Audio Spectrogram Transformer (AST) [gong2021ast]. AST represents audio as Mel-spectrograms [ustubioglu2023mel, hwang2020mel], turning temporal signals into two-dimensional time–frequency grids before applying patch-based Transformer processing with fixed-size patches ($16\times 16$). While AST reuses ImageNet-trained weights [deng2009imagenet, gong2021ast], SSAST [gong2022ssast] processes pre-training on AudioSet [gemmeke2017audio] and LibriSpeech [panayotov2015librispeech] rather than reusing ImageNet-trained weights for extending AST.
Models such as Wav2Vec 2.0 [baevski2020wav2vec] and HuBERT [devlin2019bert] process raw wavforms using 1D CNN-Transformer architectures. Wav2Vec 2.0 is composed of seven convolutional layers equipped with LayerNorm and GELU activations, and its large-scale pre-training on LibriSpeech [panayotov2015librispeech] and Libri-Light [kahn2020libri] reportedly required about 7.5 days on 128 GPUs. HuBERT follows a similar CNN-Transformer architecture to Wav2Vec 2.0 but extends it with a BERT-style [devlin2019bert] masked prediction mechanism that integrates a waveform encoder. Like Wav2Vec 2.0, AMAuT adopts a 1D CNN-Transformer structure but differs fundamentally: AMAuT trains from scratch without pre-training, uses 1D CNNs on Mel-spectrogram horizontal tokens, and replaces LayerNorm + GELU with BatchNorm + ReLU. These changes improve both efficiency and accuracy (99.82% vs. 99.4% on AudioMNIST) between Wav2Vec 2.0 and AMAuT.
The spectrogram transformers [zhang2022spectrogram] contributed a Tempo-Frequency sequential (TFS) attention for separating mel-spectrogram outputs within frequency direction instead of $16\times 16$ regions in AST [gong2021ast] that are similar to the horizontal tokenization of AMAuT (see subsection 3.3). However, [zhang2022spectrogram] inherits the methodology of AST and ViT, and then [zhang2022spectrogram] employs a different model architecture, as it incorporates 2D CNN layers.
@@ 2.2 Test-Time Domain Adaptation
Test-Time domain adaptation (TTDA) only uses the test set for performance improvement by adapting (partial) model parameters of a trained model from a source domain to a target domain [liang2024comprehensive]. In AMAuT, this is achieved through a composite loss function $\mathcal{L}_{TTDA}$ that combines three complementary objectives: Nuclear-Norm Maximization, Entropy Minimization, and Generalized Entropy (see Eq.7, Eq.8, and Eq.9 in Appendix A). These components have been applied separately in previous studies, such as CoNMix [kumar2023conmix] for test-time domain adaptation, TENT [wang2021tent] for entropy-based test-time batch adaptation, and SGEM [kim2023sgem] for Automatic Speech Recognition, but AMAuT unifies them for the first time. This integration allows efficient test-time domain adaptation and improves robustness under moderate distribution shifts without additional supervision.
@@ 3.2 Multiview Augmentation
  • Training (four augmentations): (1) random left or right time shift (0-17%), (2) Gaussian noise (noise-to-audio ratio = 0.015, and regenerate noise each time), and (3-4) random clips from two background noises (“dude miaowing” and “pick noise” from SpeechCommands [warden2018speech]) with an SNR of 50 dB.
@@ 3.5 Vertical Embedding
AMAuT introduces two learnable tokens: the class token (CLS) and the tail token (TAL) (see Eq. 3 and Figure 3). Similar to DeiT [touvron2021training], AMAuT positions CLS and TAL at opposite ends of the token sequence.
@@
This component follows the Transformer [vaswani2017attention] architecture and comprises multiple transformer blocks. Two important configurations are adopted:
@@
    The dropout rate for both multi-head attention and MLP is set to 0%, consistent with models such as SSAST [gong2022ssast], AST [gong2021ast], and CoNMix [kumar2023conmix].
@@ 3.7 Test-Time Adaptation / Augmentation ($TTA^{2}$)
This subsection investigates an uncommon application of Test-time Domain Adaptation (TTDA): improving accuracy when training and test distributions are similar. Typically, TTDA is employed to address distribution shifts, but we explore its utility even when such a shift is minimal [liang2024comprehensive].
TTDA enhances performance by adapting a trained model to the test data distribution without requiring labels [liang2024comprehensive]. During TTDA, all test features are visible, but labels remain hidden. The model updates its internal parameters to better align with the target domain before final prediction.
@@
where $\alpha$, $\beta$, and $\gamma$ are scaling coefficients (see Table12 in Appedix C).
@@
To further enhance inference robustness, AMAuT adopts test-time augmentation (TTAu) [kimura2021understanding]. Unlike TTDA, TTAu does not update model parameters; instead, it refines predictions by averaging outputs from multiple augmented versions of the same input. Three TTAu strategies are evaluated: Augmentation-driven refinement (Aug), Multi-training-based refinement (Mlt), and Hybrid refinement (Hyb) (see Eq.10, Eq.11, and Eq.12 in Appendix A, separately).
@@ 3.8 Learning Rate and Optimization
where $e$ denotes the epoch number, $10\leq\lambda\leq 40$ controls the rate of change, $200\geq\eta\geq 40$ defines the cardinality. See more details of hyperparameter settings from Table 12 in Appendix C.
@@ 4 Experiments and Results
AMAuT integrates several architectural and training strategies, including augmentation-driven multiview learning, conv1 + conv7 + conv1 CNN bottlenecks, flexible handling of arbitrary sample rates and audio lengths, learnable positional embeddings, and $TTA^{2}$. Collectively, these components enable AMAuT to achieve strong performance when trained entirely from scratch, without relying on large-scale pre-training.
@@ 4.1 Datasets
To evaluate the proposed AMAuT framework, this study employed five publicly available benchmark datasets: AudioMNIST (AM) [audiomnist2023], SpeechCommands V1 (SC1) & V2 (SC2) [warden2018speech], VocalSound (VS) [gong_vocalsound], and CochlScene (CS) [jeong2022cochlscene].
@@
4.2 Comparative Performance
Table 3 compares AMAuT with several state-of-the-art pre-trained models across the five datasets.
Table 3: Comparison of prediction accuracy (%) between the proposed AMAuT and state-of-the-art pre-trained models on five benchmark datasets.
Algorithm Set Accuracy(%)
AMAuT SC1 96.31
SSAST [gong2022ssast] SC1 96.20
AMAuT SC2 96.21
EAT [chen2024eat] SC2 98.30
AST [gong2021ast] SC2 98.11
PDC [chrysos2022augmenting] SC2 97.8
AMAuT AM 99.82
HuBERT [hsu2021hubert, la2024benchmarking] AM 99.95
XLS-R [babu2021xls, la2024benchmarking] AM 99.86
Wav2Vec2 [baevski2020wav2vec, la2024benchmarking] AM 99.42
AMAuT VS 92.26
Qwen-Audio [chu2023qwen] VS 92.89
EfficientNet [gong_vocalsound, tan2019efficientnet] VS 90.5
AMAuT CS 72.58
Audio Flamingo [kong2024audio] CS 83.0
Qwen-Audio [chu2023qwen] CS 79.5
  • 1
    Performance values for baseline models are obtained from their respective original publications.
AMAuT achieved 96.31% on SC1 and 99.82% on AM, closely matching or exceeding prior results, while maintaining minimal computational overhead.
@@ 4.3 Ablation Studies
Table 4: Comparison of test accuracies between sequential augmentation and augmentation-driven multiview learning on CS and SC1 before applying $TTA^{2}$.
@@
  • 1
    “1-Aug” and “4-Aug” indicate the use of one or four augmented views, respectively.
  • 2
    The arrows ($\uparrow$, $\downarrow$) denote performance improvement or degradation relative to the baseline.
@@
Consistency Analysis:
To further verify that augmentation-driven multiview learning improves convergence stability and generalization, a consistency analysis was conducted using the agreement rate metric (see Eq. 15 in Appendix B).
Table 5: Comparison of agreement rates between one and four augmentation-driven multiview configurations for AMAuT across five benchmark datasets before applying $TTA^{2}$.
Set Length 1-Aug 4-Aug
[…]
AM 1 s 0.9995 N/A
  • 1
    “1-Aug” and “4-Aug” indicate the use of one or four augmented views, respectively.
  • 2
    Agreement rate ranges from 0 to 1, where 1 denotes perfect consistency.
  • 3
    For training AM, multiple views are unnecessary, as a single view already achieves above 99.25% accuracy (see Table 7).
As shown in Table 5, the four-view configuration consistently yielded higher agreement rates across CS, VS, SC1, and SC2. This improvement indicates that augmentation-driven multiview learning enhances model stability and convergence reliability. The results thus provide empirical evidence that multiview augmentation effectively strengthens the robustness and reproducibility of the AMAuT framework.
@@
In many existing CNN-based audio classification architectures, relatively small convolutional kernels are commonly used. For example, HuBERT [hsu2021hubert] and Wav2Vec 2.0 [baevski2020wav2vec] employ bottleneck structures such as conv2 + conv3, while other hybrid CNN-Transformer architectures like ENACT-Heart [han2025enact] and CNN-TE [chen2024hybrid] use kernel sizes of 3, 2, and 1. These designs typically include only one or two large convolutional layers, such as conv14 or conv16, at the head or tail of the CNN stack, rather than throughout its body.
@@
Table 6: Performance comparison between conv3 and conv7 kernel structures on CS and SC1 before applying $(TTA)^{2}$.
@@
  • 1
    “Length” indicates audio length in seconds.
  • 2
    “1-Aug” / “4-Aug” denotes the number of augmented views used during testing.
  • 3
    Evaluated on a single RTX 4090 GPU
@@
Table 7: Test accuracy improvement (%) from different refinement strategies, including Aug, Mlt, Hyb, and TTDA, applied to AMAuT.
@@
  • 1
    For AudioMNIST (AM), results are reported over five folds (AM0–AM4).
@@
Table 8: Adaptation time comparison for different $TTA^{2}$ methods in AMAuT using a single RTX 4090 GPU.
@@
  • 1
    Time is shown in seconds (s) or minutes (m) per method.
  • 2
    “TTDA-Epoch” denotes time costing and the number of epochs used for performing TTDA.
@@
CLS+TAL Tokens
In AMAuT, the CLS (class) and TAL (tail) tokens are positioned at opposite ends of the token sequence (Similar to DeiT [touvron2021training]).
Table 9: Comparing different positional embedding strategies on CS and VS before applying $TTA^{2}$.
Set Length Method Accuracy (%)
CS 10 s CLS+TAL 68.87
CS 10 s CLS 67.13
VS 12 s CLS+TAL 91.06
VS 12 s CLS 90.56
  • 1
    See CLS token in [dosovitskiy2020image, touvron2021training].
  • 2
    ’Length’ denotes audio length in seconds.
The results summarized in Table 9 show that CLS+TAl tokens perform better than a single CLS token.